Solve the following system of equations: 5x-2y=8 3x-y=7 thanks!
Answer by tommyt3rd(5050)

5x-2y=8
m2: 3x-y=7

5x-2y=8
6x-2y=14

Swap them and subtract:
x=6

3(6)-y=7
y=18-7=11
y=11

:)